# Visualizing Thought: Conceptual Diagrams Enable Robust Planning in LMMs

### v1

@@ 5 Conclusion
In this paper, we introduced a zero-shot framework enabling LMMs to robustly solve combinatorial planning problems by integrating textual reasoning with self-generated conceptual diagrams. Our key contributions include: (1) autonomously generating conceptual diagrams directly from textual descriptions; (2) structuring inference as a visual graph-of-thought with multiple chains of intermediate textual and diagrammatic representations; (3) employing beam search and depth-wise backtracking for efficient combinatorial search; and (4) significantly outperforming GPT-4o and the specialized reasoning model o1-preview across multiple challenging planning domains.
Future Work.
Besides enhancing reasoning, our generated conceptual diagrams can directly benefit downstream tasks such as grounded video and image generation, where they can be used as explicit scene layouts which have shown to significantly improve quality [32, 31]. Our structured inference naturally produces accurate diagrams grounded in provided natural language constraints, offering a promising pathway toward fully automatic visual generation for complex, long-horizon tasks.
Supplementary Material

### v2

@@ 5 Conclusion
In this paper, we introduced Visual Thinking, a zero-shot framework enabling LMMs to reason using text and self-generated conceptual diagrams to solve combinatorial planning problems. Our key contributions include: (1) autonomously generating conceptual diagrams directly from textual descriptions; (2) inference as a visual Graph-of-Thought with multiple chains of intermediate textual and diagrammatic representations; (3) employing beam search and depth-wise backtracking for efficient search; and (4) significantly outperforming GPT-4o and the specialized reasoning model o1-preview across multiple challenging planning domains.
Limitations.
As with any search-based inference method, our framework incurs additional computational cost and inference time by exploring multiple chains of thought and including visual representations for each state. Moreover, while setting a larger branching factor or beam width for the search tree can improve performance, it leads to higher computational cost and latency. However, we note that our method proved more computationally efficient than prior text-only search approaches like standard Graph-of-Thought (GoT) and RAP.
Future Work and Real-World Applications.
This work demonstrates a path for LMMs to move beyond purely textual reasoning toward a more powerful, human-like process that integrates reasoning through visual abstractions. Future work can extend this framework from combinatorial planning to more abstract domains, enabling LMMs to generate multimodal outputs such as software architecture diagrams, figures to visualize scientific hypotheses and causal dependencies, or tailored visual aids for educational contexts.
Our approach also has immediate applications in real-world combinatorial problems governed by flexible, natural-language constraints that challenge traditional algorithms. For instance, it could optimize warehouse layouts or product planograms based on qualitative rules (e.g., “group seasonal items together”) or resolve complex scheduling problems involving nuanced user preferences. In these contexts, the diagrams serve a dual role: providing a structured world model for the model to reason through and an intuitive validation tool for humans. This ability to reason and communicate multimodally is crucial for building AI that can tackle complex scientific, creative, and planning tasks.